OpenStudy (anonymous):

@sinusoidal @djaydeej88 Josh bought 20 shares of Netflix at $117.98. A year later, he bought 20 more shares at $124.32. He later sold all of his shares at the price of $128.48. Josh's broker charges him 2% of each share as a transaction fee. After all of his transactions, how much money did Josh make? A. $93.50 B. $106.27 C. $112.31 D. $91.79

OpenStudy (anonymous):

you will need to break it down like this: 1. total cost at $117.98 per share 2. total cost at $124.32 per share 3. apply the 2% to #1 and #2 4. solve like what you did in the previous example.

OpenStudy (calculusxy):

20*117.98 + 20*124.32 = $4846.00 $4846.00*1.02 = $4942.92 40*128.48 = $5139.20 0.02*5139.20 = $102.78 $5139.20 - $102.78 = $5036.42 $5036.42 - $4942.92 = $93.50, So the answer should be A. yahoo answers